Find the indicated IQ score. The graph to the right depicts IQ scores of adults, and those scores are normally distributed with a mean of 100 and a standard deviation of 15 0.3446 The indicated 1Q score is (Round to the nearest whole number as needed.)
Answer #1

Solution :

mean = \mu = 100

standard deviation = \sigma = 15

Using standard normal table,

P(Z > z) = 0.3446

1 - P(Z < z) =0.3446

P(Z < z) = 1 - 0.3446 = 0.6554

P(Z < 0.40) = 0.6554

z = 0.40

Using z-score formula,

x = z * \sigma + \mu

x = 0.40 * 15 + 100 = 106

The indicated IQ score is 106 .